OpenStudy (anonymous):

Solve 6x ≡ 9 (mod 21).

OpenStudy (amistre64):

6x/21 = nR9 right?

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

reduce it all by a factor of 3 maybe ... 2x/7 = n R3 ?? i dont do mods enough to be confident in them

OpenStudy (amistre64):

9 9 +21 = 30 <- heres one 30+21 = 51 51+21 = 72 <- heres one 72+21 = 93 93+21 = 114 <- heres one seems to be a pattern

OpenStudy (amistre64):

6(5,12,19,...)

OpenStudy (anonymous):

right what amistre said. you can write \[6x\equiv 9(21)\] \[6x\equiv 30(21)\] \[x\equiv 5(21)\]

OpenStudy (amistre64):

6(5+7n) ; n>=0 maybe

myininaya (myininaya):

\[6x=9+21k\] I like to enter in values of k to see what makes x an integer k=0 => x not integer k=1 => x=5 k=2 => x not integer k=3 => x=12 k=4 => x not integer k=5 => x=19 k=6 => x not intger k=7 => x=26 so there is a pattern x=5+7i for any integer i

OpenStudy (anonymous):

oh did i screw up? maybe \[6x\equiv 9(21)\] \[2x\equiv 3(7)\] \[2x\equiv 10(7)\] \[x\equiv 5(7)\] is a better answer

OpenStudy (amistre64):

i saw that but wasnt gonna say nuthin ;)

OpenStudy (anonymous):

lol back to the dungeon for me